Short answer problems A manager has developed a payoff table that under two possible states of nature. indicates the profits assocdated with a set o Alt S1 S2 10 -2 55) If the manager uses maximin as the decision criterion, which of the alternatives should she choose? Answer: Alt Diff: 2 Page Section Heading: Dedision Making without Probabilities Keywords: maximin eriterion AACSB: Analytical thinking 15) If the manager uses minimax regret as the decision criterion, which of the alternatives would she Answer Alt S1 S2 worst 12 0 12 3 (min regret) Diff: 2 Page Ref: 540 Section Heading: Decision Making without Probabilities Keywords: minimax regret criterion AACSB: Analytical thinking 6) Use the expected value criterion to select the best altermative. Assume thathe probabilty of S s equal to 0.4 Answer. rmxD EV (Alt 3)- Answer: Diff: 2 Page Ref: 555 Section Heading: Decision Making without Probabilities Keywords expected value criterion AACSB: Analytical thinking
media%2Fb1b%2Fb1b3e8fc-7d7b-4335-a8de-8b
Answer #1

55)

Using Maximin decision criterion, that alternative is selected whose minimum payoff is the maximum of the minimum payoffs of all the alternatives.

Minimum payoff of Alt 1 = MIN(10, 2) = 2

Minimum payoff of Alt 2 = MIN(-2, 8) = -2

Minimum payoff of Alt 3 = MIN(8, 5) = 5

Maximum of the above = MAX(2, -2, 5) = 5, It pertains to Alt 3.

So according to Maximin criterion, Alt 3 should be selected

15)

Refer the regret table given in the question,

Maximum regret of Alt 1 = MAX(0, 6) = 6

Maximum regret of Alt 2 = MAX(12, 0) = 12

Maximum regret of Alt 3 = MAX(2, 3) = 3

Minimum of the above is 3, which pertains to Alt 3.

So, according to Minimax regret criterion, Alt 3 should be selected.

16)

Probability of S2 is 0.4, so probability of S1 = 1-0.4 = 0.6

EV(Alt 1) = .6*10 + .4*2 = 6.8

EV(Alt 2) = .6*(-2) + .4*8 = 2

EV(Alt 3) = .6*8 + .4*5 = 6.8

EV of Alt 1 and 3 are the highest. So,

according to Expected Value criterion, Alt 1 or 3 should be selected

17)

EVwPI = .6*Maximum payoff under S1 + .4*Maximum payoff under S2

= .6*10 + .4*8 = 9.2

EVPI = EVwPI - EVw/oPI = 9.2 - 6.8 = 2.4